# FP4DiT: Towards Effective Floating Point Quantization for Diffusion Transformers

### v2

@@ Abstract
Diffusion Models (DM) have revolutionized the text-to-image visual generation process. However, the large computational cost and model footprint of DMs hinders practical deployment, especially on edge devices. Post-training quantization (PTQ) is a lightweight method to alleviate these burdens without the need for training or fine-tuning. While recent DM PTQ methods achieve W4A8 on integer-based PTQ, two key limitations remain: First, while most existing DM PTQ methods evaluate on classical DMs like Stable Diffusion XL, 1.5 or earlier, which use convolutional U-Nets, newer Diffusion Transformer (DiT) models like the PixArt series, Hunyuan and others adopt fundamentally different transformer backbones to achieve superior image synthesis. Second, integer (INT) quantization is prevailing in DM PTQ but does not align well with the network weight and activation distribution, while Floating-Point Quantization (FPQ) is still under-investigated, yet it holds the potential to better align the weight and activation distributions in low-bit settings for DiT. In this paper, we introduce FP4DiT, a PTQ method that leverages FPQ to achieve W4A6 quantization. Specifically, we extend and generalize the Adaptive Rounding PTQ technique to adequately calibrate weight quantization for FPQ and demonstrate that DiT activations depend on input patch data, necessitating robust online activation quantization techniques. Experimental results demonstrate that FP4DiT outperforms integer-based PTQ at W4A6 and W4A8 precision and generates convincing visual content on PixArt-$\alpha$, PixArt-$\Sigma$ and Hunyuan in terms of several T2I metrics such as HPSv2 and CLIP.
@@ 1 Introduction
Nevertheless, similar to U-Nets, DiTs utilize a lengthy denoising process that incurs a high computational inference cost. One method to alleviate this burden is quantization (Dettmers et al., 2022; Yao et al., 2022), which reduces the bit-precision of neural network weights and activations. As the first Post-Training Quantization (PTQ) schemes for DMs, PTQ4DM (Shang et al., 2023b) and Q-Diffusion (Li et al., 2023) demonstrate that the range and distribution of U-Net activations crucially depend on the diffusion timestep. More recent state-of-the-art works like TFMQ-DM (Huang et al., 2023) specialize quantization for U-Net timestep conditioning which may not generalize to newer DiTs. Further, methods like ViDiT-Q (Zhao et al., 2024) adapt outlier suppression technique (Xiao et al., 2023) to DiTs, but overlook broader advantages of prior DM PTQ like weight reconstruction (Nagel et al., 2020; Li et al., 2021).
Moreover, the prevailing datatypes in existing DM PTQ literature (Shang et al., 2023b; Li et al., 2023; Huang et al., 2023; Zhao et al., 2024; So et al., 2024; Feng et al., 2025; Mills et al., 2025) are integer-based (INT), which provide uniformly distributed values (Nahshan et al., 2021) unlike the non-uniform distribution of weights and activations in modern neural networks (Shen et al., 2024). Thus, PTQ for text-to-image (T2I) DiTs below W4A8 precision (4-bit weights and 8-bit activations) without severely compromising generation quality remains an open challenge.
In this paper, we present FP4DiT, which achieves W4A6 PTQ on Diffusion Transformers with non-uniform Floating-Point Quantization (FPQ) (Kuzmin et al., 2022), thus achieving high quantitative and qualitative T2I performance. Besides, by introducing FPQ, FP4DiT not only aligns the quantization levels better with the weight and activation distribution with negligible computational overhead, it also massively reduces the cost of weight calibration by over 8$\times$. Our detailed contributions are summarized as follows:
@@
We apply FP4DiT as a PTQ method on T2I DiT models, namely PixArt-$\alpha$, PixArt-$\Sigma$, and Hunyuan. To verify the effectiveness of FP4DiT, we conduct extensive experiments on T2I tasks such as the Human Preference Score v2 (HPSv2) benchmark  (Wu et al., 2023a) and MS-COCO dataset (Lin et al., 2014b), to outperform existing methods like Q-Diffusion (Li et al., 2023), TFMQ-DM (Huang et al., 2023), ViDiT-Q (Zhao et al., 2024) and Q-DiT (Chen et al., 2025) at the W4A8 and W4A6 precision levels. Additionally, we perform a human preference study which demonstrates the superiority of FP4DiT-generated images.
@@ 2 Related Work
Diffusion Transformers (DiT) (Peebles & Xie, 2023) replace the classical convolutional U-Net (Rombach et al., 2022) backbone with a modified Vision Transformer (ViT) (Dosovitskiy et al., 2020) to increase scalability. Although the introduction of DiT architectures in newer DMs (Chen et al., 2024b; a; Esser et al., 2024; Li et al., 2024; Labs, ; Xie et al., 2024) enables the generation of high-quality visual content (Brooks et al., 2024), DiTs still suffer from a computationally expensive diffusion process, rendering deployment on edge devices impractical and cumbersome. Further, addressing this weakness for DiTs poses unique challenges compared to U-Nets, and is a focus of this work.
Quantization is a neural network compression technique that involves reducing the bit-precision of weights and activations to lower hardware metrics like model size, inference latency and memory consumption (Nagel et al., 2021). The objective of quantization research is to reduce bit-precision as much as possible while preserving overall model performance (Ma et al., 2024). There are two main classes of quantization: Quantization-Aware-Training (QAT) (Sui et al., 2025; He et al., 2023; Feng et al., 2025) and Post-Training Quantization (PTQ) (Li et al., 2021; Mills et al., 2025). Specifically, PTQ is more lightweight and neither requires re-training nor substantial amounts of data. Rather, PTQ requires a small amount of data to calibrate quantization scales (Nagel et al., 2020), typically in a block-wise manner (Li et al., 2021). However, while most PTQ methods rely on uniformly-distributed integer (INT) quantization techniques (Krishnamoorthi, 2018; Jacob et al., 2018), recent literature highlights the advantages of low-bit floating point quantization (FPQ) (Wu et al., 2023b; Liu et al., 2023) for Large Language Models (LLM). Therefore, in this paper we investigate the challenges in applying FPQ to DM PTQ.
@@
In contrast, LLM quantization (Dettmers et al., 2022; Frantar et al., 2022) operates on generative AI transformers. The key challenge is that multi-billion parameter transformers tend to generate outlier hidden states that are difficult to effectively quantize while preserving end-to-end performance (Lee et al., 2024a; Shang et al., 2023a). This can be addressed by leveraging the learnable affine shift of layernorm operations to adjust transformer attention and feedforward weights (Xiao et al., 2023; Lin et al., 2024b; a). However, DiTs use Adaptive Layernorm (AdaLN) (Perez et al., 2018), which ties the affine shift to the timestep embedding, so these methods are less applicable. Additionally, LLM quantization typically features more lightweight calibration (Ashkboos et al., 2024) as parameter-heavy models make advanced PTQ (Nagel et al., 2020; Li et al., 2021) costly. However, DiTs typically have fewer parameters and benefit from diverse calibration sets to cover multiple timesteps. Thus, FP4DiT leverages prior work on PTQ calibration to quantize DiTs.
Finally, some early research exists on DiT PTQ (Chen et al., 2025). HQ-DiT (Liu & Zhang, 2024) apply FPQ to class-conditional ImageNet (Deng et al., 2009) DiTs, but do not consider text-to-image (T2I) models like the PixArt (Chen et al., 2024b; a) series. ViDiT-Q (Zhao et al., 2024) utilizes fine-grained techniques including channel balancing, mixed-precision and LLM outlier suppression, it does not incorporate weight reconstruction (Nagel et al., 2020; Li et al., 2021) from prior DM PTQ works. In contrast, this work aggregates knowledge from existing DM PTQ methods and refines them for application on T2I DiT models.
@@ 3 Methodology
Figure 2: T2I DiT block diagram. In PixArt-$\alpha$/$\Sigma$, all DiT blocks share the same AdaLN-single MLP for time conditions. The scale and shift for layer normalization in DiT blocks depend on the embedding from AdaLN-single and the layer-specific training embedding. Colored blocks demarcate quantizable weight layers from activations and normalizations.
[…]
@@
The key advantage of FPQ, especially at low-bit precision for quantization, is that they enjoy a richer granularity of value distributions owing to the numerous ways we can vary the allocation of exponent and mantissa bits. This is analogous to the introduction of the ‘BFloat16’ (Kalamkar et al., 2019) format, which achieves superiority over the older IEEE standard 754 ‘Float16’ (Kahan, 1996) in certain deep machine algorithms (Lee et al., 2024b) by allocating 8-bits towards the exponent, as the larger ‘Float32’ format does. Broadly, an $n$-bit floating point datatype posses $n-1$ possible distributions as $m\in[0,n-1]$, and even adopts the uniform distribution of the corresponding $n$-bit integer format when $m=n-1$.
@@ 3.2 AdaRound for FP
where $min$ and $max$ denotes the quantization threshold. $h(\mathbf{V})$ is the rectified sigmoid function proposed by (Louizos et al., 2017):
@@ 4.1 Experimental Settings
PixArt-$\alpha$ Full Precision W16A16 32.56 31.06 30.76 29.67 31.01 34.05 0.3075
Q-Diffusion W4A8 24.18 23.43 22.91 22.15 23.17 52.20 0.3017
TFMQ-DM W4A8 27.88 26.03 25.14 24.91 25.99 64.73 0.3066
ViDiT-Q W4A8 17.93 17.35 16.81 17.59 17.42 39.11 0.2900
Q-DiT W4A8 28.49 26.91 27.55 26.56 27.38 36.17 0.3062
FP4DiT (ours) W4A8 28.63 26.79 27.59 26.72 27.43 30.37 0.3076
Q-Diffusion W4A6 12.63 13.39 13.32 10.65 12.50 70.96 0.2868
TFMQ-DM W4A6 24.02 23.36 22.72 23.04 23.29 90.09 0.3015
ViDiT-Q W4A6 16.71 16.28 16.23 16.56 16.44 56.54 0.2827
Q-DiT W4A6 24.66 22.59 23.29 23.04 23.40 43.91 0.3015
FP4DiT (ours) W4A6 24.57 23.08 23.30 23.26 23.55 40.61 0.3031
PixArt-$\Sigma$ Full Precision W16A16 33.07 31.58 31.54 30.49 31.67 36.94 0.3139
Q-Diffusion W4A8 27.30 26.11 26.06 25.06 26.13 36.89 0.3050
TFMQ-DM W4A8 24.95 23.59 23.19 22.73 23.62 62.98 0.2975
ViDiT-Q W4A8 27.10 26.45 26.24 24.49 26.07 37.17 0.2562
Q-DiT W4A8 27.74 26.23 26.05 24.67 26.17 32.03 0.3048
FP4DiT (ours) W4A8 27.95 26.29 26.16 24.67 26.27 31.58 0.3064
Q-Diffusion W4A6 24.07 22.42 22.47 21.72 22.67 74.83 0.3027
TFMQ-DM W4A6 19.30 18.46 18.85 17.50 18.53 154.13 0.2346
ViDiT-Q W4A6 24.84 23.23 23.48 21.94 23.37 87.47 0.2425
Q-DiT W4A6 25.67 23.57 23.60 22.64 23.87 73.20 0.2894
FP4DiT (ours) W4A6 26.91 25.55 25.22 23.93 25.40 42.21 0.3040
Hunyuan Full Precision W16A16 33.72 31.84 31.52 31.24 32.08 59.08 0.3102
Q-Diffusion W4A8 26.00 24.74 24.84 24.26 24.96 82.43 0.3006
TFMQ-DM W4A8 28.77 27.63 27.67 26.15 27.56 89.39 0.3075
ViDiT-Q W4A8 28.74 26.79 26.49 27.20 27.31 82.22 0.3099
Q-DiT W4A8 28.91 27.27 27.15 26.72 27.52 85.33 0.3088
FP4DiT (ours) W4A8 28.96 27.75 27.47 26.94 27.78 81.94 0.3102
Q-Diffusion W4A6 14.90 14.83 15.32 14.24 14.83 255.69 0.2277
TFMQ-DM W4A6 13.73 13.31 13.16 14.51 13.68 258.02 0.2520
ViDiT-Q W4A6 15.11 15.20 15.41 14.17 14.97 265.52 0.2559
Q-DiT W4A6 13.41 13.15 13.48 12.94 13.24 262.31 0.2312
FP4DiT (ours) W4A6 15.23 14.94 15.57 14.27 15.00 255.06 0.2562
Table 2: Quantitative evaluation results for PixArt-$\alpha$, PixArt-$\Sigma$ and Hunyuan in terms of HPSv2, FID, and CLIP score. Specifically, for each configuration, we generate 10k images (5k for Hunyuan) using COCO 2014 validation set prompts. Best and second best results in bold and italics, respectively.

### v3

@@ Abstract
Diffusion Models (DM) have revolutionized the text-to-image visual generation process. However, the large computational cost and model footprint of DMs hinders practical deployment, especially on edge devices. Post-training quantization (PTQ) is a lightweight method to alleviate these burdens without the need for training or fine-tuning. While recent DM PTQ methods achieve W4A8 (i.e., 4-bit weights and 8-bit activations) on integer-based PTQ, two key limitations remain: First, while most existing DM PTQ methods evaluate on classical DMs like Stable Diffusion XL, 1.5 or earlier, which use convolutional U-Nets, newer Diffusion Transformer (DiT) models like the PixArt series, Hunyuan and others adopt fundamentally different transformer backbones to achieve superior image synthesis. Second, integer (INT) quantization is prevailing in DM PTQ but does not align well with the network weight and activation distribution, while Floating-Point Quantization (FPQ) is still under-investigated, yet it holds the potential to better align the weight and activation distributions in low-bit settings for DiT. In this paper, we introduce FP4DiT, a PTQ method that leverages FPQ to achieve W4A6 quantization. Specifically, we extend and generalize the Adaptive Rounding PTQ technique to adequately calibrate weight quantization for FPQ and demonstrate that DiT activations depend on input patch data, necessitating robust online activation quantization techniques. Experimental results demonstrate that FP4DiT achieves higher CLIP, ImageReward and HPSv2 performance compared to integer-based PTQ at the W4A6 and W4A8 precision levels while generating convincing visual content on $\alpha, PixArt$-$\Sigma$ and Hunyuan. Code is available at https://github.com/cccrrrccc/FP4DiT.
@@ 1 Introduction
Nevertheless, similar to U-Nets, DiTs utilize a lengthy denoising process that incurs a high computational inference cost. Approaches to reducing this cost generally aim to reduce the inference cost of the denoising network (Fang et al., 2023; Castells et al., 2024) or reduce the number of denoising steps (Yin et al., 2024; Salimans & Ho, 2022). One specific method to reduce the denoiser cost is quantization (Dettmers et al., 2022; Yao et al., 2022), which reduces the bit-precision of neural network weights and activations. As the first Post-Training Quantization (PTQ) schemes for DMs, PTQ4DM (Shang et al., 2023b) and Q-Diffusion (Li et al., 2023) demonstrate that the range and distribution of U-Net activations crucially depend on the diffusion timestep. More recent state-of-the-art works like TFMQ-DM (Huang et al., 2024) specialize quantization for U-Net timestep conditioning which may not generalize to newer DiTs. Further, methods like ViDiT-Q (Zhao et al., 2024) adapt outlier suppression technique (Xiao et al., 2023) to DiTs, but overlook broader advantages of prior DM PTQ like weight reconstruction (Nagel et al., 2020; Li et al., 2021).
Moreover, the prevailing datatypes in existing DM PTQ literature (Shang et al., 2023b; Li et al., 2023; Huang et al., 2024; Zhao et al., 2024; So et al., 2024; Feng et al., 2025; Mills et al., 2025) are integer-based (INT), which provide uniformly distributed values (Nahshan et al., 2021) unlike the non-uniform distribution of weights and activations in modern neural networks (Shen et al., 2024). Thus, PTQ for text-to-image (T2I) DiTs below W4A8 precision (4-bit weights and 8-bit activations) without severely compromising generation quality remains an open challenge.
[…]
@@
We apply FP4DiT as a PTQ method on T2I DiT models, namely $\alpha, PixArt$-$\Sigma$, and Hunyuan. To verify the effectiveness of FP4DiT, we conduct experiments on the MS-COCO dataset (Lin et al., 2014b) and the Human Preference Score v2 (HPSv2) T2I benchmark, to outperform existing methods like Q-Diffusion (Li et al., 2023), TFMQ-DM (Huang et al., 2024), ViDiT-Q (Zhao et al., 2024) and Q-DiT (Chen et al., 2025a) at the W4A8 and W4A6 precision levels. Additionally, we perform a human preference study which demonstrates the superiority of FP4DiT-generated images.
@@ 2 Related Work
Diffusion Transformers (DiT) (Peebles & Xie, 2023) replace the classical convolutional U-Net (Rombach et al., 2022) backbone with a modified Vision Transformer (ViT) (Dosovitskiy et al., 2020) to increase scalability. Although the introduction of DiT architectures in newer DMs (Chen et al., 2024b; a; Esser et al., 2024; Li et al., 2024; Labs, ; Xie et al., 2024) enables the generation of high-quality visual content (Brooks et al., 2024), DiTs still suffer from a computationally expensive diffusion process, rendering deployment on edge devices impractical and cumbersome. Methods to reduce this cost might aim to compress the denoiser neural network by pruning (Fang et al., 2023; Castells et al., 2024), sparsifying key operations (Chen et al., 2025b), or by targeting a reduction in the number of denoising steps needed to generate an image (Yin et al., 2024; Salimans & Ho, 2022). This work focuses on the former objective while also aiming to discover some of the challenges associated with achieving DiT improvement compared to prior U-Nets.
Quantization is a neural network compression technique that involves reducing the bit-precision of weights and activations to lower hardware metrics like model size, inference latency and memory consumption (Nagel et al., 2021). The objective of quantization research is to reduce bit-precision as much as possible while preserving overall model performance (Ma et al., 2024). There are two main classes of quantization: Quantization-Aware-Training (QAT) (Sui et al., 2025; He et al., 2023; Feng et al., 2025) and Post-Training Quantization (PTQ) (Li et al., 2021; Mills et al., 2025; Frantar & Alistarh, 2022). Specifically, PTQ is more lightweight and neither requires re-training nor substantial amounts of data. Rather, PTQ requires a small amount of data to calibrate quantization scales (Nagel et al., 2020), typically in a block-wise manner (Li et al., 2021). However, while most PTQ methods rely on uniformly-distributed integer (INT) quantization techniques (Krishnamoorthi, 2018; Jacob et al., 2018; Zhang et al., 2025), recent literature highlights the advantages of low-bit floating point quantization (FPQ) (Wu et al., 2023b; Liu et al., 2023) for Large Language Models (LLM). Therefore, in this paper we investigate the challenges in applying FPQ to DM PTQ.
@@
In contrast, LLM quantization (Dettmers et al., 2022; Frantar et al., 2022; Chee et al., 2023; Liu et al., 2024; Zhang et al., 2024) operates on generative AI transformers. The key challenge is that multi-billion parameter transformers tend to generate outlier hidden states that are difficult to effectively quantize while preserving end-to-end performance (Lee et al., 2024a; Shang et al., 2023a). This can be addressed by leveraging the learnable affine shift of layernorm operations to adjust transformer attention and feedforward weights (Xiao et al., 2023; Lin et al., 2024b; a). However, DiTs use Adaptive Layernorm (AdaLN) (Perez et al., 2018), which ties the affine shift to the timestep embedding, so these methods are less applicable. Additionally, LLM quantization typically features more lightweight calibration (Ashkboos et al., 2024) as parameter-heavy models make advanced PTQ (Nagel et al., 2020; Li et al., 2021) costly. However, DiTs typically have fewer parameters and benefit from diverse calibration sets to cover multiple timesteps. Thus, FP4DiT leverages prior work on PTQ calibration to quantize DiTs.
Finally, some early research exists on DiT PTQ (Chen et al., 2025a). HQ-DiT (Liu & Zhang, 2024) apply FPQ to class-conditional ImageNet (Deng et al., 2009) DiTs, but do not consider text-to-image (T2I) models like the PixArt (Chen et al., 2024b; a) series. ViDiT-Q (Zhao et al., 2024) utilizes fine-grained techniques including channel balancing, mixed-precision and LLM outlier suppression, it does not incorporate weight reconstruction (Nagel et al., 2020; Li et al., 2021) from prior DM PTQ works. In contrast, this work aggregates knowledge from existing DM PTQ methods and refines them for application on T2I DiT models.
@@ 3 Methodology
where $s$ is scale, $z$ is the zero point, and $\left\lfloor\cdot\right\rceil$ is operation rounding-to-nearest. INT quantization rounds the values to a range with $2^{n}$ points. Specifically, the range is always a uniform grid, whose size decreases by half each time $n$ decreases by 1.
@@
Figure 2: T2I DiT block diagram. In PixArt-$\alpha$$\Sigma$, all DiT blocks share the same AdaLN-single MLP for time conditions. The scale and shift for layer normalization in DiT blocks depend on the embedding from AdaLN-single and the layer-specific training embedding. Colored blocks demarcate quantizable weight layers from activations and normalizations.
@@
The key advantage of FPQ, especially at low-bit precision for quantization, is that they enjoy a richer granularity of value distributions owing to the numerous ways we can vary the allocation of exponent and mantissa bits. This is analogous to the introduction of the ‘BFloat16’ (Kalamkar et al., 2019) format, which achieves superiority over the older IEEE standard 754 ‘Float16’ (Kahan, 1996) in certain deep machine algorithms (Lee et al., 2024b) by allocating 8-bits towards the exponent, as the larger ‘Float32’ format does. Broadly, an $n$-bit floating point datatype posses n1n-1 possible distributions as $m\in[0,n-1]$, and even adopts the uniform distribution of the corresponding $n$-bit integer format when m=n1m=n-1.
@@ 3.2 AdaRound for FP
where minmin and maxmax denotes the quantization threshold. $h(\mathbf{V})$ is the rectified sigmoid function proposed by (Louizos et al., 2017):
@@
(a) AdaRound for Integers
(b) AdaRound Applied to Floats
(c) Scale-Aware AdaRound
@@
(a) DiT Input Value Box Plot
(b) Output Scale of 7th DiT Block
(c) 7th DiT Block Output Box Plot
@@ 4.1 Experimental Settings
PixArt-$\alpha$ Full Precision W16A16 32.56 31.06 30.76 29.67 31.01 34.05 0.3075 0.7037
Q-Diffusion W4A8 24.18 23.43 22.91 22.15 23.17 52.20 0.3017 0.1855
TFMQ-DM W4A8 27.88 26.03 25.14 24.91 25.99 64.73 0.3066 0.4706
ViDiT-Q W4A8 17.93 17.35 16.81 17.59 17.42 39.11 0.2900 0.4716
Q-DiT W4A8 28.49 26.91 27.55 26.56 27.38 36.17 0.3062 0.4979
FP4DiT (ours) W4A8 28.63 26.79 27.59 26.72 27.43 30.37 0.3076 0.5509
Q-Diffusion W4A6 12.63 13.39 13.32 10.65 12.50 70.96 0.2868 -0.1825
TFMQ-DM W4A6 24.02 23.36 22.72 23.04 23.29 90.09 0.3015 0.1299
ViDiT-Q W4A6 16.71 16.28 16.23 16.56 16.44 56.54 0.2827 0.1296
Q-DiT W4A6 24.66 22.59 23.29 23.04 23.40 43.91 0.3015 0.1323
FP4DiT (ours) W4A6 24.57 23.08 23.30 23.26 23.55 40.61 0.3031 0.1353
PixArt-$\Sigma$ Full Precision W16A16 33.07 31.58 31.54 30.49 31.67 36.94 0.3139 0.9223
Q-Diffusion W4A8 27.30 26.11 26.06 25.06 26.13 36.89 0.3050 0.2285
TFMQ-DM W4A8 24.95 23.59 23.19 22.73 23.62 62.98 0.2975 0.1047
ViDiT-Q W4A8 27.10 26.45 26.24 24.49 26.07 37.17 0.2562 0.2832
Q-DiT W4A8 27.74 26.23 26.05 24.67 26.17 32.03 0.3048 0.2679
FP4DiT (ours) W4A8 27.95 26.29 26.16 24.67 26.27 31.58 0.3064 0.2927
Q-Diffusion W4A6 24.07 22.42 22.47 21.72 22.67 74.83 0.3027 0.0828
TFMQ-DM W4A6 19.30 18.46 18.85 17.50 18.53 154.13 0.2346 -0.2281
ViDiT-Q W4A6 24.84 23.23 23.48 21.94 23.37 87.47 0.2425 -0.1156
Q-DiT W4A6 25.67 23.57 23.60 22.64 23.87 73.20 0.2894 0.1614
FP4DiT (ours) W4A6 26.91 25.55 25.22 23.93 25.40 42.21 0.3040 0.2085
Hunyuan Full Precision W16A16 33.72 31.84 31.52 31.24 32.08 59.08 0.3102 0.8701
Q-Diffusion W4A8 26.00 24.74 24.84 24.26 24.96 82.43 0.3006 0.6874
TFMQ-DM W4A8 28.77 27.63 27.67 26.15 27.56 89.39 0.3075 0.6873
ViDiT-Q W4A8 28.74 26.79 26.49 27.20 27.31 82.22 0.3099 0.7423
Q-DiT W4A8 28.91 27.27 27.15 26.72 27.52 85.33 0.3088 0.7539
FP4DiT (ours) W4A8 28.96 27.75 27.47 26.94 27.78 81.94 0.3102 0.7669
Q-Diffusion W4A6 14.90 14.83 15.32 14.24 14.83 255.69 0.2277 -1.9236
TFMQ-DM W4A6 13.73 13.31 13.16 14.51 13.68 258.02 0.2520 -1.9007
ViDiT-Q W4A6 15.11 15.20 15.41 14.17 14.97 265.52 0.2559 -1.8123
Q-DiT W4A6 13.41 13.15 13.48 12.94 13.24 262.31 0.2312 -1.1822
FP4DiT (ours) W4A6 15.23 14.94 15.57 14.27 15.00 255.06 0.2562 -1.7880
Table 2: Quantitative evaluation results for $\alpha, PixArt$-$\Sigma$ and Hunyuan in terms of HPSv2, FID, CLIP and ImageReward (IR) score. Specifically, for each configuration, we generate 10k images (5k for Hunyuan) using COCO 2014 validation set prompts. Best and second best results in bold and italics, respectively.